OpenStudy (anonymous):

What is the slop (-5,5) and (2,3)

OpenStudy (y2o2):

(y-y1)/(x-x1)

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (3-5)/(2-(-5)) m = (3-5)/(2+5) m = -2/7 So the slope of the line that passes through the two points is -2/7